OpenStudy (katherinena13):

Before I was old enough to drive, my brothers had to take turns driving me around. In exchange for this, my mom told them that she would pay for their car insurance. The base insurance rate was $2500 minus $50 for every year the driver is insured with the company. My oldest brother, Jeff, had earned a good driver discount of $500 off. However, he also drove a sports car, which added $250 to his premium. Greg had not been driving long enough to earn a discount, but he drove a little pick-up so there was no additional cost. (cont.)

OpenStudy (katherinena13):

The cost of their insurance premiums can be described using the functions below: Jeff's insurance: J(x) = 2500 - 50x - 500 + 250 Greg's insurance: G(x) = 2500 - 50x where x is the amount of years the driver is insured with the company. 4. In the fifth year, how much more did my mom have to pay for Greg than for Jeff? Use f(x)=2x-5 and g(x)=12-x^2 to evaluate the expression

OpenStudy (katherinena13):

@jojokiw3

OpenStudy (katherinena13):

@TheSmartOne @Hero @nincompoop